# Toward quantum-safe scalable networks: an open, standards-aware key management framework

### v1

@@ V Discussion and Conclusions
This work has presented a complete architecture and mechanism for E2E quantum key establishment across QKD networks, combining centralized path computation with decentralized key relay execution. By abstracting the key management complexity through a virtual KMS layer and relying on an SDN-based centralized quantum security controller, the proposed solution enables seamless key delivery for applications over both direct and multi-hop trusted relay paths.
The system has been designed to ensure interoperability and specifications compliance. It builds on the ETSI GS QKD 014 API for key request and delivery, ETSI GS QKD 015 for SDN-based control, and anticipates integration with the emerging ETSI GS QKD 020 for KMS-to-KMS key relay.
The use of a centralized controller (QuSeC) for KMS discovery and relay path computation brings notable advantages in efficiency and scalability. With a global view of the network, the QuSeC is able to compute optimal relay paths, based on different configurable criteria such as hop count or available keyrate, and install them in all involved KMSs. Moreover, the QuSeC maintains session state, allowing subsequent key requests by the peer application to proceed without requiring installation of the relay path and avoiding redundant configuration in trusted relay scenarios.
Regarding performance, the proposed mechanism requires querying the QuSeC for every key request, even for simple direct-path requests. While this guarantees consistent operation across all scenarios, it may introduce overhead and latency in environments where paths are stable and known. This limitation could be addressed through caching mechanisms at the vKMS level, where the results of recent path computations can be reused for repeated requests, thus reducing latency and controller load.
In summary, the proposed architecture provides a practical, standards-aligned solution for E2E key delivery across QKD infrastructures. It supports a broad range of topologies, simplifies application integration, and remains extensible for future developments in KMS interoperability and controller intelligence.

### v2

@@ V Discussion and Conclusions
This work has presented a complete architecture and mechanism for E2E quantum key establishment across QKD networks, combining centralized path computation with decentralized key relay execution. By abstracting the key management complexity through a virtual KMS layer and introducing an SDN-based quantum security controller, the proposal enables seamless key delivery over both direct and multi-hop trusted relay paths.
The design ensures interoperability and compliance with ETSI specifications, building on the 014 API for key delivery, 015 for SDN-based control, and anticipating integration with the emerging 020 for KMS-to-KMS key relay. The centralized QuSeC efficiently computes relay paths based on configurable metrics such as hop count or available key rate and maintains session state to avoid redundant configurations.
While this proposal follows a logically centralized model, distributed control approaches could enhance resilience by avoiding a single point of failure. However, QuSeC centralization provides global visibility and simplified interoperability with existing standards. This approach also remains scalable under high-load conditions with frequent key requests, since control-plane traffic for KMS discovery and path setup is orders of magnitude lower than data traffic. In large-scale deployments, QuSeC replication and state synchronization across redundant instances (using east/westbound interfaces defined for centralized controllers) could provide load balancing and high availability, mitigating performance bottlenecks and avoiding single-point failures.
Regarding performance, experimental results from demonstrate the feasibility of the proposal, showing E2E key establishment time scales efficiently with the number of hops. Overall, the results demonstrate that centralized control introduces negligible overhead and that caching at the vKMS level could further reduce latency and controller load for repeated path computations.
Overall, the proposed framework provides a practical, standards-aligned solution for E2E key delivery across QKD infrastructures. It supports a broad range of topologies, simplifies application integration, and remains extensible for future developments in KMS interoperability and controller intelligence.
@@ Acknowledgments
VI Biography Section
Ane Sanz received her BSc degree in Telecommunication Engineering in 2019 and her MSc degree in Telecommunication Engineering in 2021 from the University of the Basque Country (UPV/EHU). She joined the Communications Engineering Department of the UPV/ EHU as a researcher in the I2T Research Lab (Engineering and Research on Telematics) in 2018. Her research interests include quantum communications applied to security and networking. Currently, she is a PhD student in Mobile Network Information and Communication Technologies at the UPV/EHU.
Asier Atutxa received his BSc and MSc degrees in Telecommunication Engineering in 2018 and 2020, and his PhD in Mobile Network Information and Communication Technologies from the University of the Basque Country (UPV/EHU) in 2025. He joined the Communications Engineering Department of the UPV/EHU as a researcher in the I2T Research Lab (Engineering and Research on Telematics) in 2018, where he currently works. His research interests include cybersecurity, software-defined networks, and virtualization applied to industrial environments.
David Franco received the B.Sc. and M.Sc. degrees in Telecommunication Engineering from the University of the Basque Country (UPV/EHU) in 2016 and 2018, respectively. He also received the Ph.D. in Mobile Network Information and Communication Technologies from the UPV/EHU in 2025. He joined the UPV/EHU in 2016 as a researcher in the I2T (Engineering and Research on Telematics) research group, and he has been a PhD Assistant Professor in the Communications Engineering Department of UPV/EHU since 2025. His research focuses on software-defined networking and network function virtualisation applied to traffic engineering and secure communication systems.
Jasone Astorga received the B.Sc. and M.Sc. degrees in Telecommunication Engineering and the Ph.D. degree from the University of the Basque Country (UPV/EHU), in 2004 and 2013, respectively. From 2004 to 2007, she worked with Nextel S.A., a telecommunications enterprise. She joined the UPV/EHU in 2007 as a lecturer and a researcher with the I2T Research Laboratory. She is currently an Assistant Professor with the UPV/EHU. She has led and contributed to numerous local, national, and European research projects in cybersecurity for industrial environments, Software Defined Networking (SDN), Network Function Virtualization (NFV), and IoT security. She has co-authored numerous journal papers and conference contributions, and has supervised four PhD candidates and numerous master’s and bachelor projects.
Eduardo Jacob {Senior Member, IEEE} received the M.S. in Industrial Engineering and Ph.D. degree in Information and Communication Technologies in 2001 from the University of the Basque Country (UPV/EHU), Spain. He is currently Full Professor with the Department of Communications Engineering at UPV/EHU. He is the I2T Research Laboratory principal investigator and SmartQuanT4E and SmartNets4E research infrastructures director. His research interests include computer networks, network security, and advanced network architectures, with a focus on quantum safe and time-sensitive networking. He has authored and co-authored numerous publications in international journals and conferences and actively participates and lead national and European research projects.
Diego Lopez received the M.S. in Physics from the University of Granadada, and the Ph.D. in Physics from the University of Seville in 2000. He joined Telefonica in 2011 as a Senior Technology Expert, where he is currently in charge of the Technology Exploration activities within the GCTIO Unit. Before joining Telefónica he spent some years in the academic sector, dedicated to research on network services, and was appointed member of the High-Level Expert Group on Scientific Data Infrastructures by the European Commission. Diego is currently focused on applied research in network infrastructures, with a special emphasis on data-driven management, new architectures, security, and quantum communications. Diego is an ETSI Fellow and chairs the ETSI TC DATA and ISG ZSM.